OpenStudy (anonymous):

What is the value of x when -3x+7x=-12

OpenStudy (dan815):

x(-3+7)=-12

OpenStudy (anonymous):

so???

OpenStudy (dan815):

simplify

OpenStudy (dan815):

7 appmes minus 3 apples is what?

OpenStudy (anonymous):

oh umm...

OpenStudy (igreen):

-3 + 7 = ? @!Call~Me~Addi!

OpenStudy (dan815):

for 4 apples u are giving the shop owner 12 dollars, so u are losing 12 dollars 4 Apples = -12 how much money are u losing for just 1 apple then?

OpenStudy (igreen):

Can you add that? @!Call~Me~Addi!

OpenStudy (anonymous):

12/4 haha @dan815

OpenStudy (dan815):

ok but remember thats how much the shop owner is getting, what he gets u are losing so its -12/4

OpenStudy (anonymous):

what shop owner?

OpenStudy (anonymous):

You're confusing the poor kid. @dan815 LOL

OpenStudy (anonymous):

You are making this kinda of hard

OpenStudy (igreen):

Lol @Linchpin

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

okay are we going to get an answer?

OpenStudy (anonymous):

\[\href{ http://openstudy.com/code-of-conduct }{\text{Code of Conduct}}\]

OpenStudy (dan815):

tryna help us not be scared of x and equal and minus signs

OpenStudy (anonymous):

Linch, stop posting nonsense

OpenStudy (anonymous):

@Owlfred Not working. ._.

OpenStudy (anonymous):

what Dan?

OpenStudy (anonymous):

@PaulaLovesSchool13 Rude...

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

I said not to post nonsense, it's spamming

OpenStudy (anonymous):

sorry !! had to help my little sister so let me read everything you guys said ...

OpenStudy (anonymous):

uh.....

OpenStudy (igreen):

Just add -3 + 7..what do you get? @!Call~Me~Addi!

OpenStudy (anonymous):

@PaulaLovesSchool13 I'm not spamming..I just wanted to link @dan815 to the kok,thx.

OpenStudy (anonymous):

Hmmm...next time send it in a message privately

OpenStudy (anonymous):

wow... you guys talk alot and @iGreen one sec

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

-3x+7x=-12 4x=-12 x=-12/4 @!Call~Me~Addi! Getting it?

OpenStudy (anonymous):

Had a feeling you would say that

OpenStudy (anonymous):

Linch, catch *throws medal*

OpenStudy (igreen):

Haha :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

*falls down* Yeah Thanks. :D @PaulaLovesSchool13

OpenStudy (anonymous):

Was the medal too heavy? o.o

OpenStudy (anonymous):

I wasn't ready for it...\(\LARGE\color{pink}{\ddot\smile}\)

OpenStudy (anonymous):

lol anyway this person is not answering so I am gonna go, see ya

OpenStudy (anonymous):

sorry gtg